# Ihara’s Lemma for $\operatorname{GL}_{d}$: the limit case

### v4

@@ 1.2. Generalisations of Ihara’s Lemma
The first natural one developed by Clozel, Harris and Taylor in their first proof of Sato-Tate theorem [CHT08], focuses on the $H^{0}$ with coefficients in ${\mathbb{F}}_{l}$ of a zero dimensional Shimura variety associated to higher dimensional definite division algebras. More precisely consider a totally real field $F^{+}$ and a imaginary quadratic extension $E/{\mathbb{Q}}$ and define $F=F^{+}E$. We then consider $\overline{G}/{\mathbb{Q}}$ an unitary group with $\overline{G}({\mathbb{Q}})$ compact so that $\overline{G}$ becomes an inner form of $\operatorname{GL}_{d}$ over $F$. This means, cf. §2.3, we have fixed a division algebra $\overline{B}$ with center $F$, of dimension $d^{2}$, provided with an involution of the second kind such that its restriction to $F$ is the complex conjugation. We moreover suppose that at every place $w$ of $F$, either $\overline{B}_{w}$ is split or a local division algebra.
@@
To a maximal ideal $\mathfrak{m}$ of ${\mathbb{T}}(\overline{K}^{v})$ is associated a Galois $\overline{\mathbb{F}}_{l}$-representation $\overline{\rho}_{\mathfrak{m}}$, cf. §4.2. We consider the case where this representation is irreducible. Note in particular that such an $\mathfrak{m}$ is then not pseudo-Eisenstein in the usual terminology.
@@ 1.3. Main result
The basic idea111this explains the hypothesis on the existence of $p_{0}$ in the statement, cf. §4.1, as in [Boy22], is to introduce geometry and move from the Shimura variety associated to $\overline{G}$ which is of dimension zero, to another Shimura variety $\operatorname{Sh}_{K}$ associated to some reductive group $G$ and level $K$, of strictly positive dimension, so that $S_{\overline{G}}(\overline{K},\overline{\mathbb{F}}_{l})$ appears in a certain cohomology group of some sheaf over $\operatorname{Sh}_{K}$. The strategy is then to construct a filtration, coming from geometry, of this cohomology group so that the graded parts, which are expected to be more easy to handle with, also verify the genericity property of there irreducible sub-spaces.
More explicitly we study the middle degree cohomology group of the KHT Shimura variety $\operatorname{Sh}_{K^{v}(\infty)}$ associated to some similitude group $G/{\mathbb{Q}}$ such that $G({\mathbb{A}}_{\mathbb{Q}}^{\infty,p})\cong\overline{G}({\mathbb{A}}_{\mathbb{Q}}^{\infty,p})$, cf. §2.3 for more details, and with level $K^{v}(\infty):=\overline{K}^{v}$ meaning finite level outside $v$ and infinite level at $v$. The localization at $\mathfrak{m}$ of the cohomology groups of $\operatorname{Sh}_{K^{v}(\infty)}$ can be computed as the cohomology of the geometric special fiber $\operatorname{Sh}_{K^{v}(\infty),\bar{s}_{v}}$ of $\operatorname{Sh}_{K^{v}(\infty)}$, with coefficient in the complex of nearby cycles $\Psi_{K^{v}(\infty),v}$.
The Newton stratification of $\operatorname{Sh}_{K^{v}(\infty),\bar{s}_{v}}$ gives us a filtration of $\Psi_{K^{v}(\infty),v}$, cf. [Boy20], and so a filtration $\operatorname{Fil}^{\bullet}(K^{v}(\infty))$ of $H^{d-1}(\operatorname{Sh}_{K^{v}(\infty),\bar{\eta}_{v}},\overline{\mathbb{Z}}_{l})_{\mathfrak{m}}$ and the main point of [Boy22] is to prove that the modulo $l$ reduction of each graded part of this filtration verifies the Ihara property, i.e. each of their irreducible sub-space are generic. To realize this strategy we need first the cohomology groups of $\operatorname{Sh}_{K^{v}(\infty)}$ to be torsion free: this point is now essentially settled by the main result of [Boy23a]. More crucially the previous filtration $\operatorname{Fil}^{\bullet}(K^{v}(\infty))$ should be strict, i.e. its graded parts have to be torsion free, cf. theorem 3.2.3.
@@
One problem about this last point is that the ¯$\overline{\mathbb{Q}}_{l}$-cohomology is split. For any irreducible automorphic representation $\Pi$ of $G({\mathbb{A}})$ cohomological for, say, the trivial coefficients, the ¯$\overline{\mathbb{Z}}_{l}$-cohomology defines a lattice $\Gamma(\Pi)$ of $(\Pi^{\infty})^{K^{v}(\infty)}\otimes\sigma(\Pi)_{v}$ whose modulo $l$ reduction gives a subspace of the $\overline{\mathbb{F}}_{l}$-cohomology: Ihara’s lemma predicts that the socle of this subspace is still generic, i.e. it gives informations about the lattice $\Gamma(\Pi)$. We then see that non splitness of $\operatorname{Fil}^{\bullet}(K^{v}(\infty))\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{F}}_{l}$ should be understood in a very flexible point of view.
One possible strategy is, using the fact that the ¯$\overline{\mathbb{Q}}_{l}$-cohomology is split, to start from the filtration $\operatorname{Fil}^{\bullet}(K^{v}(\infty))$ and modify it in order to arrive to another one where all the modulo $l$ reduction of the graded parts fulfill the Ihara property i.e. their irreducible subspaces are generic. The main ingredient to construct modifications of filtrations is to consider following situations:
@@
  • We can then define $\widetilde{\operatorname{gr}^{k-1}}$ and $\widetilde{\operatorname{gr}^{k}}$ with
@@
    obtained by taking $\widetilde{\operatorname{gr}^{k}}:=X\cap(\operatorname{gr}^{k}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{Q}}_{l})$, and where $T$ is torsion. Passing modulo $l$, we then obtained a priori two distinct filtrations.
  • We can define a $\overline{\mathbb{F}}_{l}$-monodromy operator for the Galois action at the place $v$ [^2] Note that over $\overline{\mathbb{F}}_{l}$ the usual arithmetic approach for defining the nilpotent monodromy operator, is hopeless because, up to consider a finite extension of $F_{v}$, such a $\overline{\mathbb{F}}_{l}$-representation has a trivial action of the inertia group. We are looking for a geometric monodromy operator $N^{geo}_{v}$ which then exists whatever are the coefficients, ¯$\overline{\mathbb{Q}}_{l}$, ¯$\overline{\mathbb{Z}}_{l}$ and $\overline{\mathbb{F}}_{l}$, compatible with tensor products. One classical construction is known in the semi-stable reduction case, cf. [Ill94] §3, which corresponds to the case where the level at $v$ of our Shimura variety is of Iwahori type. [^3] This corresponds to automorphic representations $\Pi$ such that the cuspidal support of $\Pi_{v}$ is made of unramified characters, and so with the weak form of Ihara’s lemma of definition 1.2.3. Using our knowledge of the ¯$\overline{\mathbb{Z}}_{l}$-nearby cycles described completely in [Boy23b], we can construct such a geometric nilpotent monodromy operator which generalizes the semi-stable case, cf. §3.3.
  • Taking this geometric monodromy operator, we then obtain a cohomological monodromy operator $N_{v,\mathfrak{m}}^{coho}$ acting on $H^{0}(\operatorname{Sh}_{K,\bar{s}_{v}},\Psi_{K^{v}(\infty),v})_{\mathfrak{m}}$ One of the main point, cf theorem 3.2.3, is that the graded parts of the filtration of $H^{0}(\operatorname{Sh}_{K,\bar{s}_{v}},\Psi_{K^{v}(\infty),v})_{\mathfrak{m}}$ induced by the Newton filtration on the nearby cycles spectral sequence, are all torsion free, so that in particular we are in position to understand quite enough the action of $\overline{N}^{coho}_{v,\mathfrak{m}}:=N^{coho}_{v,\mathfrak{m}}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{F}}_{l}$ on $H^{0}(\operatorname{Sh}_{K,\bar{s}_{v}},\Psi_{K^{v}(\infty),v})_{\mathfrak{m}}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{F}}_{l}$, and prove that its nilpotency order is as large as possible.
  • Note that as $\overline{\rho}_{\mathfrak{m}}$ is supposed to be irreducible, then the modulo $l$ reduction of the monodromy operator acting on $\rho_{\widetilde{\mathfrak{m}}}$ does not depend on the choice of the prime ideal $\widetilde{\mathfrak{m}}\subseteq\mathfrak{m}$ so that it is usually trivial.
Finally, as $N_{v}^{coho}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{F}}_{l}$ is far from being trivial, there should be non split extensions between the graded parts of $\operatorname{Fil}^{\bullet}(K^{v}(\infty))$.
However this strategy seems difficult to implement directly because of counting problems: to deal with finite number of representations you need to work with a finite level at the place $v$ and then pass to the limit. It seems first difficult to count liftings of a fixed representation and secondly when increasing the level, it should be not easy to glue back things together. Our approach in some sense consists to consider all the liftings together using typicness of the cohomology, cf. §4.3. The proof finally goes into three main steps:
  • we first prove, cf. theorem 3.2.3, that the filtration of the middle cohomology group constructed from the filtration of stratification of the nearby cycles perverse sheaf, has torsion free graded parts, otherwise the all cohomology would have non trivial torsion classes which is not the case by [Boy23a];
  • secondly, §3.3, using results from [Boy23b] about various filtrations of stratification of the nearby cycle perverse sheaf over ¯$\overline{\mathbb{Z}}_{l}$, we define an integral monodromy operator on the nearby cycles perverse sheaf whose order of nilpotency modulo $l$ coincides with those over ¯$\overline{\mathbb{Q}}_{l}$;
  • this integral geometric monodromy operator gives us a monodromy operator on $\rho_{\mathfrak{m}}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{F}}_{l}$, a representation with coefficients in an artinian local ring defined as the modulo $l$ reduction of the image of some Hecke algebra acting on the middle cohomology group with finite level. Working at various finite level and using Matlis duality for artinian ring, we prove, §4.3, the genericity of the socle of the middle cohomology group at infinite level at $v$, for the action of $\operatorname{GL}_{d}(F_{v})$.
@@ 2. Preliminaries
2.1. Representations of $\operatorname{GL}_{d}(L)$
Consider a finite extension $L/{\mathbb{Q}}_{p}$ with residue field ${\mathbb{F}}_{q}$. We denote by |||-|| - | its absolute value. For a representation $\pi$ of $\operatorname{GL}_{d}(L)$ and $n\in\frac{1}{2}{\mathbb{Z}}$, set
@@ 2.1. Representations of $\operatorname{GL}_{d}(L)$
For $\pi_{1}$ and $\pi_{2}$ representations of respectively $\operatorname{GL}_{n_{1}}(L)$ and $\operatorname{GL}_{n_{2}}(L)$, we will denote by
$\pi_{1}\times\pi_{2}:=\operatorname{ind}_{P_{n_{1},n_{1}+n_{2}}(L)}^{\operatorname{GL}_{n_{1}+n_{2}}(L)}\pi_{1}\{\frac{n_{2}}{2}\}\otimes\pi_{2}\{-\frac{n_{1}}{2}\},$
@@
Recall that a representation $\varrho$ of $\operatorname{GL}_{d}(L)$ is called cuspidal (resp. supercuspidal) if it is not a subspace (resp. subquotient) of a proper parabolic induced representation. When the field of coefficients is of characteristic zero, these two notions coincides, but this is no more true over $\overline{\mathbb{F}}_{l}$.
@@
(see [Zel80] §9 and [Boy10] §1.4) Let $g$ be a divisor of $d=sg$ and $\pi$ an irreducible cuspidal ¯$\overline{\mathbb{Q}}_{l}$-representation of $\operatorname{GL}_{g}(L)$. The induced representation
@@
For any $\underline{s}\in\operatorname{Dec}(s)$, we the denote by $\mathrm{st}_{\underline{s}}(\pi)$ the associated irreducible sub-quotient of (1). Following Zelevinsky, we fix this bijection such that $\operatorname{Speh}_{s}(\pi)$ corresponds to $(s)$ and $\mathrm{st}_{s}(\pi)$ to $(1,\cdots,1)$. The Lubin-Tate representation $LT_{h,s}(\pi)$ will also appear in the following, it corresponds with $(\overbrace{1,\cdots,1}^{h},s-h)$.
@@
(cf. [Vig96] III.5.10) Let $\pi$ be an irreducible cuspidal representation of $\operatorname{GL}_{g}(K)$ with a stable $\overline{\mathbb{Z}}$-lattice [^4] We say that $\pi$ is integral., then its modulo $l$ reduction is irreducible and cuspidal (but not necessary supercuspidal).
@@
  • the modulo $l$ reduction of every irreducible cuspidal representation of $\operatorname{GL}_{g}(L)$ for $g\leq d$, is supercuspidal [^5] In the banal case this is not always the case but it is when the cuspidal support contains only characters.: with the notation of [Boy11] proposition 1.3.5, $m(\varrho)=l>d$ for any irreducible $\overline{\mathbb{F}}_{l}$-supercuspidal representation $\varrho$.
  • For a $\overline{\mathbb{F}}_{l}$-irreducible supercuspidal representation $\varrho$ of $\operatorname{GL}_{g}(L)$, the parabolic induced representation $\varrho\times\cdots\times\varrho$, with $s$ copies of $\varrho$, is semi-simple with irreducible constituants the modulo $l$ reduction of the set of elements of $\{\mathrm{st}_{\underline{s}}(\pi)\hbox{ such that }\underline{s}\in\operatorname{Dec}(s)\}$, where $\pi$ is any cuspidal representation whose modulo $l$ reduction is isomorphic to $\varrho$.
Concerning the notion of genericity, consider the mirabolic subgroup $M_{d}(L)$ of $\operatorname{GL}_{d}(L)$ as the set of matrices with last row $(0,\cdots,0,1)$: we denote by
$V_{d}(L)=\{(m_{i,j})\in M_{d}(L):~{}m_{i,j}=\delta_{i,j}\hbox{ for }j<d\}.$
its unipotent radical. We fix a non trivial character $\psi$ of $L$ and let $\theta$ be the character of $V_{d}(L)$ defined by $\theta((m_{i,j}))=\psi(m_{d-1,d})$. For $G=\operatorname{GL}_{r}(L)$ or $M_{r}(L)$, we denote by $\operatorname{alg}(G)$ the abelian category of smooth representations of $G$ and, following [BZ77], we introduce
$\Psi^{-}:\operatorname{alg}(M_{d}(L))\longrightarrow\operatorname{alg}(\operatorname{GL}_{d-1}(L)),$
@@
$\Phi^{-}:\operatorname{alg}(M_{d}(L))\longrightarrow\operatorname{alg}(M_{d-1}(L)),$
defined by $\Psi^{-}=r_{V_{d},1}$ (resp. $\Phi^{-}=r_{V_{d},\theta}$) the functor of $V_{d}$ coinvariants (resp. $(V_{d},\theta)$-coinvariants), cf. [BZ77] 1.8. For $\tau\in\operatorname{alg}(M_{d}(L))$, the representation
@@
is called the $k$-th derivative of $\tau$. If $\tau^{(k)}\neq 0$ and $\tau^{(m)}=0$ for all $m>k$, then $\tau^{(k)}$ is called the highest derivative of $\tau$. In the particular case where $k=d$, there is an unique irreducible representation $\tau_{nd}$ of $M_{d}(L)$ with derivative of order $d$.

### v5

@@ 1.2. Generalisations of Ihara’s Lemma
The first natural one developed by Clozel, Harris and Taylor in their first proof of the Sato-Tate theorem [CHT08], focuses on the $H^{0}$ with coefficients in ${\mathbb{F}}_{l}$ of a zero dimensional Shimura variety associated to higher dimensional definite division algebras. More precisely consider a totally real field $F^{+}$ and a imaginary quadratic extension $E/{\mathbb{Q}}$ and define $F=F^{+}E$. We then consider $\overline{G}/{\mathbb{Q}}$ an unitary group with $\overline{G}({\mathbb{Q}})$ compact so that $\overline{G}$ becomes an inner form of $\operatorname{GL}_{d}$ over $F$. This means, cf. §2.3, we have fixed a division algebra $\overline{B}$ with center $F$, of dimension $d^{2}$, provided with an involution of the second kind such that its restriction to $F$ is the complex conjugation. We moreover suppose that at every place $w$ of $F$, either $\overline{B}_{w}$ is split or a local division algebra.
@@
To a maximal ideal $\mathfrak{m}$ of ${\mathbb{T}}(\overline{K}^{v})$ is associated a Galois $\overline{\mathbb{F}}_{l}$-representation $\overline{\rho}_{\mathfrak{m}}$, cf. §5.2. We consider the case where this representation is irreducible. Note in particular that such an $\mathfrak{m}$ is then non pseudo-Eisenstein in the usual terminology.
@@ 1.3. Main results
The basic idea111this explains the hypothesis on the existence of $p_{0}$ in the statement, cf. §5.1, as in [Boy22], is to introduce geometry and move from the Shimura variety associated to $\overline{G}$ which is of dimension zero, to another Shimura variety $\operatorname{Sh}_{K}$ associated to some reductive group $G$ and level $K$, of strictly positive dimension, so that $S_{\overline{G}}(\overline{K},\overline{\mathbb{F}}_{l})$ appears in a certain cohomology group of some sheaf over $\operatorname{Sh}_{K}$. The first point in our strategy is to work in characteristic zero and then consider modulo $l$ reduction. For this we will need to fix the following data.
Notation 1.3.3.
Consider a coefficient field $L$ which is a large enough finite extension of ${\mathbb{Q}}_{l}$, with ring of integers ${\mathcal{O}}_{L}$ and residue field ${\mathcal{O}}_{L}/\varpi_{F}{\mathcal{O}}_{L}{\mathbb{F}}_{L}$ for some fixed uniformizer $\varpi_{L}$.
We then want to construct a filtration of the middle cohomology group of $\operatorname{Sh}_{K}$ with coefficients in ${\mathcal{O}}_{L}$ so that the graded parts, which are expected to be more easy to handle with, all verify the genericity property of their irreducible sub-spaces. More explicitly we study the middle degree cohomology group with coefficients in ¯$\overline{\mathbb{Z}}_{l}$, of the KHT Shimura variety $\operatorname{Sh}_{K^{v}(\infty)}$ associated to some similitude group $G/{\mathbb{Q}}$ such that $G({\mathbb{A}}_{\mathbb{Q}}^{\infty,p})\cong\overline{G}({\mathbb{A}}_{\mathbb{Q}}^{\infty,p})$, cf. §2.3 for more details, and with level $K^{v}(\infty):=\overline{K}^{v}$ meaning finite level outside $v$ and infinite level at $v$. The localization at $\mathfrak{m}$ of the cohomology groups of $\operatorname{Sh}_{K^{v}(\infty)}$ can be computed as the cohomology of the geometric special fiber $\operatorname{Sh}_{K^{v}(\infty),\bar{s}_{v}}$ of $\operatorname{Sh}_{K^{v}(\infty)}$, with coefficient in the complex of nearby cycles $\Psi_{K^{v}(\infty),v}$.
The Newton stratification of $\operatorname{Sh}_{K^{v}(\infty),\bar{s}_{v}}$ gives us a filtration of $\Psi_{K^{v}(\infty),v}$, cf. [Boy20], and so a filtration $\operatorname{Fil}^{\bullet}(K^{v}(\infty))$ of $H^{d-1}(\operatorname{Sh}_{K^{v}(\infty),\bar{\eta}_{v}},\overline{\mathbb{Z}}_{l})_{\mathfrak{m}}$ and the main point of [Boy22] is to prove that the modulo $l$ reduction of each graded part of this filtration verifies the Ihara property, i.e. each of their irreducible sub-space are generic. To realize this strategy
  • we need first the cohomology groups of $\operatorname{Sh}_{K^{v}(\infty)}$ to be torsion free: this point is now essentially settled by the main result of [Boy23a].
  • More crucially the previous filtration $\operatorname{Fil}^{\bullet}(K^{v}(\infty))$ should be strict, i.e. its graded parts have to be torsion free, cf. theorem 3.2.3.
  • For each of the graded parts, the ¯$\overline{\mathbb{Q}}_{l}$-cohomology can be described by a set of automorphic representations, each of them giving an automorphic contribution. We choose any numbering of this discrete set and for each $n$, we choose a coefficient field $L$ large enough to be able to separate the first $n$ automorphic contributions. The cohomology over ${\mathcal{O}}_{L}$ gives us a lattice which then allows us to deal with the modulo $l$ reduction meaning reduction modulo $\varpi_{L}{\mathcal{O}}_{L}$.
Remark. In the following we will just write modulo $l$ reduction for this construction.
@@
One problem about this last point is that the ¯$\overline{\mathbb{Q}}_{l}$-cohomology is split. For any irreducible automorphic representation $\Pi$ of $G({\mathbb{A}})$ cohomological for, say, the trivial coefficients, the ¯$\overline{\mathbb{Z}}_{l}$-cohomology defines a lattice $\Gamma(\Pi)$ of $(\Pi^{\infty})^{K^{v}(\infty)}\otimes\sigma(\Pi)_{v}$ whose modulo $l$ reduction gives a subspace of the $\overline{\mathbb{F}}_{l}$-cohomology: Ihara’s lemma predicts that the socle of this subspace is still generic, i.e. it gives informations about the lattice $\Gamma(\Pi)$. We then see that non splitness of $\operatorname{Fil}^{\bullet}(K^{v}(\infty))\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{F}}_{l}$ should be understood in a very flexible point of view. We then naturally face to prove the following result.
Proposition 1.3.4.
(cf. 5.5.1) The contribution $\Gamma(\Pi)$ of an automorphic representation $\Pi$ to the cohomology viewed as a subrepresentation, defines a stable lattice of $\Pi_{v}$ uniquely defined by the property that the socle of its modulo $l$ reduction is irreducible and generic.
In particular this lattice should depend only on $\Pi_{v}$ and not on the global representation $\Pi$. This statement looks similar to the Breuil lattice conjecture which is stated when $l=p$ and $K_{v}$-types. We then first prove a simple version of this result, cf. proposition 4.3.2 which can be stated as follows.
  • Consider some fixed $K_{v}$-type $\sigma_{v,\overline{\mathbb{Q}}_{l}}$
  • and a system $\lambda$ of Hecke eigenvalues associated to some automorphic representation $\Pi$ as above
  • such that $\sigma_{v,\overline{\mathbb{Q}}_{l}}$ appears with multiplicity one in $\Pi_{v}$.
Then the lattice of $\sigma_{v,\overline{\mathbb{Q}}_{l}}\cap\Gamma(\Pi)$ depends only on the modulo $l$ reduction of $\lambda$. One combinatorial problem is then to recover the information on the $\operatorname{GL}_{d}(F_{v})$-lattices from this vague obversation, cf. §5.3.2. The idea to do so is to start from the filtration $\operatorname{Fil}^{\bullet}(K^{v}(\infty))$ coming from a filtration of the nearby cycles perverse sheaf and modify it step by step, cf. §5.3, until we arrive at an automorphic filtration, cf. §5.3.1, i.e. where the graded parts correspond to the contribution of some automorphic representation. The main ingredient to construct modifications of filtrations is to consider situations as illustrated in the figure 1.
@@
  • We can then define $\widetilde{\operatorname{gr}^{k}}:=X\cap(\operatorname{gr}^{k}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{Q}}_{l})$ and the quotient $\widetilde{\operatorname{gr}^{k-1}}$: note that $T$ is torsion. We choose a $L$ finite over ${\mathbb{Q}}_{l}$, so that this diagram is defined over ${\mathcal{O}}_{L}$ and which allows to look modulo $l$, meaning modulo $\varpi_{L}{\mathcal{O}}_{L}$, where we then obtained a priori two distinct filtrations.
(a) We can define a $\overline{\mathbb{F}}_{l}$-monodromy operator for the Galois action at the place $v$ [^2] Note that over $\overline{\mathbb{F}}_{l}$ the usual arithmetic approach for defining the nilpotent monodromy operator, is hopeless because, up to consider a finite extension of $F_{v}$, such a $\overline{\mathbb{F}}_{l}$-representation has a trivial action of the inertia group. We are looking for a geometric monodromy operator $N^{geo}_{v}$ which then exists whatever are the coefficients, ¯$\overline{\mathbb{Q}}_{l}$, ¯$\overline{\mathbb{Z}}_{l}$ and $\overline{\mathbb{F}}_{l}$, compatible with tensor products. One classical construction is known in the semi-stable reduction case, cf. [Ill94] §3, which corresponds to the case where the level at $v$ of our Shimura variety is of Iwahori type. [^3] This corresponds to automorphic representations $\Pi$ such that the cuspidal support of $\Pi_{v}$ is made of unramified characters, and so with the weak form of Ihara’s lemma of definition 1.2.3. Using our knowledge of the ¯$\overline{\mathbb{Z}}_{l}$-nearby cycles described completely in [Boy23b], we can construct such a geometric nilpotent monodromy operator which generalizes the semi-stable case, cf. §3.3.
(b) Taking this geometric monodromy operator, we then obtain a cohomological monodromy operator $N_{v,\mathfrak{m}}^{coho}$ acting on $H^{0}(\operatorname{Sh}_{K,\bar{s}_{v}},\Psi_{K^{v}(\infty),v})_{\mathfrak{m}}$ One of the main point, cf theorem 3.2.3, is that the graded parts of the filtration of $H^{0}(\operatorname{Sh}_{K,\bar{s}_{v}},\Psi_{K^{v}(\infty),v})_{\mathfrak{m}}$ induced by the Newton filtration on the nearby cycles spectral sequence, are all torsion free, so that in particular we are in position to understand quite enough the action of $\overline{N}^{coho}_{v,\mathfrak{m}}:=N^{coho}_{v,\mathfrak{m}}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{F}}_{l}$ on $H^{0}(\operatorname{Sh}_{K,\bar{s}_{v}},\Psi_{K^{v}(\infty),v})_{\mathfrak{m}}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{F}}_{l}$, and prove that its nilpotency order is as large as possible.
(c) Note that as $\overline{\rho}_{\mathfrak{m}}$ is supposed to be irreducible, then the modulo $l$ reduction of the monodromy operator acting on $\rho_{\widetilde{\mathfrak{m}}}$ does not depend on the choice of the prime ideal $\widetilde{\mathfrak{m}}\subseteq\mathfrak{m}$ so that it is usually trivial.
@@
Figure 1. Exchange process
Finally, as $N_{v}^{coho}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{F}}_{l}$ is far from being trivial, there should be non split extensions between the graded parts of The heart of our proof is then divided in three main steps:
  • prove that the lattice of $\Pi_{v}$ induced by $\Gamma(\Pi)$ depends only on the modulo $l$ reduction of the system of Hecke eigenvalues of a globalization $\Pi$ of $\Pi_{v}$;
  • using the modulo $l$ monodromy operator $\overline{N}_{v}$, prove that this lattice verifies the Ihara property;
  • prove that the graded parts of our final filtration, which do not verify Ihara’s property can not give a non generic subspace of the all cohomology.
@@ 2. Preliminaries
2.1. Representations of $\operatorname{GL}_{d}(M)$
Consider a finite extension $M/{\mathbb{Q}}_{p}$ with residue field ${\mathbb{F}}_{q}$. We denote by |||-|| - | its absolute value. For a representation $\pi$ of $\operatorname{GL}_{d}(M)$ and $n\in\frac{1}{2}{\mathbb{Z}}$, set
@@ 2.1. Representations of $\operatorname{GL}_{d}(M)$
For $\pi_{1}$ and $\pi_{2}$ representations of respectively $\operatorname{GL}_{n_{1}}(M)$ and $\operatorname{GL}_{n_{2}}(M)$, we will denote by
$\pi_{1}\times\pi_{2}:=\operatorname{ind}_{P_{n_{1},n_{1}+n_{2}}(M)}^{\operatorname{GL}_{n_{1}+n_{2}}(M)}\pi_{1}\{\frac{n_{2}}{2}\}\otimes\pi_{2}\{-\frac{n_{1}}{2}\},$
@@
Recall that a representation $\varrho$ of $\operatorname{GL}_{d}(M)$ is called cuspidal (resp. supercuspidal) if it is not a subspace (resp. subquotient) of a proper parabolic induced representation. When the field of coefficients is of characteristic zero, these two notions coincides, but this is no more true over $\overline{\mathbb{F}}_{l}$.
@@
(see [Zel80] §9 and [Boy10] §1.4) Let $g$ be a divisor of $d=sg$ and $\pi$ an irreducible cuspidal ¯$\overline{\mathbb{Q}}_{l}$-representation of $\operatorname{GL}_{g}(M)$. The induced representation
@@
For any $\underline{s}\in\operatorname{Dec}(s)$, we then denote by $\mathrm{st}_{\underline{s}}(\pi)$ the associated irreducible sub-quotient of (1). Following Zelevinsky, we fix this bijection such that $\operatorname{Speh}_{s}(\pi)$ corresponds to $(s)$ and $\mathrm{st}_{s}(\pi)$ to $(1,\cdots,1)$. The Lubin-Tate representation $LT_{h,s}(\pi)$ will also appear in the following, it corresponds with $(\overbrace{1,\cdots,1}^{h},s-h)$.
Definition 2.1.3.
For $11\leq s-$ 1, we say that $\underline{s}\in\operatorname{Dec}(s)$ is $t$-small if the number of consecutive $1$ is less than $t$. We say that an irreducible subquotient of (1) is $(\pi,t)$-small if its parameter $\underline{s}\in\operatorname{Dec}(s)$ is $t$-small.
@@
(cf. [Vig96] III.5.10) Let $\pi$ be an irreducible cuspidal representation of $\operatorname{GL}_{g}(M)$ with a stable ¯$\overline{\mathbb{Z}}_{l}$-lattice [^4] As before, we fix $L/{\mathbb{Q}}_{l}$ large enough and consider an ${\mathcal{O}}_{L}$-lattice: we say that $\pi$ is integral., then its modulo $l$ reduction is irreducible and cuspidal (but not necessary supercuspidal).
@@
  • the modulo $l$ reduction of every irreducible cuspidal representation of $\operatorname{GL}_{g}(M)$ for $g\leq d$, is supercuspidal [^5] In the banal case this is not always the case but it is when the cuspidal support contains only characters.: with the notation of [Boy11] proposition 1.3.5, $m(\varrho)=l>d$ for any irreducible $\overline{\mathbb{F}}_{l}$-supercuspidal representation $\varrho$.
  • For a $\overline{\mathbb{F}}_{l}$-irreducible supercuspidal representation $\varrho$ of $\operatorname{GL}_{g}(M)$, the parabolic induced representation $\varrho\times\cdots\times\varrho$, with $s$ copies of $\varrho$, is semi-simple with irreducible constituants666some of them might be isomorphic to each others the modulo $l$ reduction of the set of elements of $\{\mathrm{st}_{\underline{s}}(\pi)\hbox{ such that }\underline{s}\in\operatorname{Dec}(s)\}$, where $\pi$ is any cuspidal representation whose modulo $l$ reduction is isomorphic to $\varrho$.
Concerning the notion of genericity, consider the mirabolic subgroup ${\mathbb{M}}_{d}(M)$ of $\operatorname{GL}_{d}(M)$ as the set of matrices with last row $(0,\cdots,0,1)$: we denote by
$\mathbb{V}_{d}(M)=\{(m_{i,j})\in{\mathbb{M}}_{d}(M):~{}m_{i,j}=\delta_{i,j}\hbox{ for }j<d\}.$
its unipotent radical. We fix a non trivial character $\psi$ of $L$ and let $\theta$ be the character of $\mathbb{V}_{d}(M)$ defined by $\theta((m_{i,j}))=\psi(m_{d-1,d})$. For $G=\operatorname{GL}_{r}(M)$ or ${\mathbb{M}}_{r}(M)$, we denote by $\operatorname{alg}(G)$ the abelian category of smooth representations of $G$ and, following [BZ77], we introduce
$\Psi^{-}:\operatorname{alg}({\mathbb{M}}_{d}(M))\longrightarrow\operatorname{alg}(\operatorname{GL}_{d-1}(M)),$
@@
$\Phi^{-}:\operatorname{alg}({\mathbb{M}}_{d}(M))\longrightarrow\operatorname{alg}({\mathbb{M}}_{d-1}(M)),$
defined by $\Psi^{-}=r_{\mathbb{V}_{d},1}$ (resp. $\Phi^{-}=r_{\mathbb{V}_{d},\theta}$) the functor of $\mathbb{V}_{d}$ coinvariants (resp. $(\mathbb{V}_{d},\theta)$-coinvariants), cf. [BZ77] 1.8. For $\tau\in\operatorname{alg}({\mathbb{M}}_{d}(M))$, the representation
@@
is called the $k$-th derivative of $\tau$. If $\tau^{(k)}\neq 0$ and $\tau^{(m)}=0$ for all $m>k$, then $\tau^{(k)}$ is called the highest derivative of $\tau$. In the particular case where $k=d$, there is an unique irreducible representation $\tau_{nd}$ of ${\mathbb{M}}_{d}(M)$ with derivative of order $d$.